Circular No. 8093

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003bv, 2003bw, 2003bx, 2003by, 2003bz, 2003ca Six apparent supernovae have been reported from CCD images — SN

2003bv by T. Boles (cf. IAUC 8089), SN 2003bw by B. Beutler and W. Li (LOTOSS/KAIT; cf. IAUC 8092), and the rest by A. Becker and the Deep Lens Survey Transient Team (cf. IAUC 7804, plus D. Loomba and R. Wilcox) from the Cerro Tololo and Kitt Peak 4-m telescopes (+ MOSAIC). The magnitudes are unfiltered unless otherwise noted.

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003bv Mar. 12.00 7^{h}22^{m}11.^{s}92 +40^{o}28′01′′.5 17.7 3′′.4 E, 5′′.3 S
2003bw Mar. 12.5 14 57 22.44 −19 12 45.9 18.0 10′′.6 E, 3′′.3 N
2003bx Mar. 3.0 5 25 04.00 −49 36 12.1 20.8V 1′′.9 W, 1′′.0 S
2003by Mar. 7.3 9 22 30.01 +29 07 19.1 22.1R 0′′.3 E, 0′′.6 N
2003bz Mar. 8.1 9 22 34.77 +30 04 42.5 23.8B 0′′.4 W, 4′′.4 S
2003ca Mar. 8.1 9 23 58.91 +30 02 39.5 23.2B 0′′.1 E, 0′′.5 N

Additional magnitudes from the respective discoverers: SN 2003bv in MCG +07-15-15, Feb. 1 UT, [19.0; 18, [19.5; Mar. 13.918, 16.8. SN 2003bw in IC 1077, 2002 July 4.3, [19.0; Mar. 3.5, (hint under poor conditions); 13.5, 18.0 (M. Schwartz). SN 2003bx, 2002 Jan. 11, V = [25; 15, B = [25; 2003 Mar. 2.0, B = 21.9. SN 2003by, 2001 Dec. 19, R = [25.5; 2003 Jan. 5, V = [25; Mar. 8.4, V = 22.7; Mar. 10.3, V = 23.0. SN 2003bz, 2001 Nov. 15, B = [25; 2003 Jan. 7, V = [25; Mar. 8.2, V = 22.3; Mar. 9.4, B = 23.9. SN 2003ca, 2001 Nov. 15, B = [25; 2003 Jan. 7, V = [25; Mar. 8.2, V = 22.2; Mar. 9.4, B = 23.4. SN 2003bv is also absent from Palomar Sky Survey red and blue plates.

UNKNOWN VARIABLE Becker also reports the discovery of a variable object located at α

= 9^{h}22^{m}38^{s}.29, δ = +29^{o}17′19′′.1, which is not near a visible host galaxy. Available magnitudes: 2001 Dec. 19 UT, R = [25.5; 2003 Jan. 5, V = [25; Mar. 7.3, R = 23.4; 8.4, V = 23.1; 10.3, V = 22.9.

COMET 66P/DU TOIT J. V. Scotti has recovered this comet with the 1.8-m Spacewatch II tele-

scope at mag 20.3–20.7 on Mar. 10 and 11. The astrometry, revised orbital elements, and an ephemeris appear on MPEC 2003-E57. The indicated correction to the prediction on MPC 40670 is ΔT = –0.25 day.

2003 March 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green